three resistors of values 2 ohms, 6 ohms, and 12 ohms respectively are connected scross a DC voltage source. what is the total current flowing this circuit? (this is all thats given)
Answer #1

Three given resistors are -

R1 = 2.0

R2 = 6.0

R3 = 12.0

The three resistors are connected across the DC voltage source. This means the three resistors are connected in parallel.

Also suppose, the voltage of the DC source = E

Equivalent resistance, R = (2.0 x 6.0 x 12.0) / (2.0x6.0 + 6.0x12.0 + 12.0x2.0) = 144.0 / (12.0 + 72.0 + 24.0)

= 144.0 / 108.0 = 1.33

Therefore, the total current through the circuit = (E / 1.33) Amp.

[Note: Put the value of the DC voltage source given in your text book.]